# Transitivity of the pure Hurwitz classes of quadratic post-critically finite polynomials

### v1

@@ 3. Portrait functions
Let $q$ be such that $f^{q}(C_{1})=C_{2}$. The branched cover $f\phi$ maps $f^{k-1}(C_{1})$ to $C_{2}$. Because $\phi$ can be defined to be supported on a neighborhood of an arc between $f^{k-1}(C_{1})$ and $\diamond$ that is disjoint from the other critical and post-critical points, $(f\phi)^{\ell}(C_{1})=f^{\ell}(C_{1})$ for $21\leq\ell\leq k-$ 2. In particular, the pre-period of $C_{1}$ under $f\phi$ consists of $\{f(C_{1}),\cdots,f^{q-1}(C_{1})\}$ and the cycle of $C_{1}$ under $f\phi$ consists of $\{C_{2}=f^{q}(C_{1}),f^{q+1}(C_{1}),\cdots,f^{k-1}(C_{1})\}$. From this, we see that the only portrait feature which remains unchanged by $f\phi$ is the number of portrait components of $\Gamma_{f}$. ∎
Function 11 (Create a new pre-period: two components).
Take as input a marked branched cover $f$ with marked set $M$ and portrait $\Gamma$ such that $\Gamma$ has two components, $\Gamma_{1}$ and $\Gamma_{2}$. Suppose further that $\Gamma_{j}$ contains $C_{j}$ and is a cycle of length at least 3. Let $\diamond$ be a preimage of $f^{2}(C_{j})$, and let $\phi$ be the half twist that swaps $\diamond$ with $f^{k-1}(C_{j})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$, and the mapping class $\phi$. Then $C_{j}$ is (strictly) pre-periodic under $C_{j}$ with pre-period $\{f(C_{j})\}$ and cycle $\{f^{2}(C_{j}),\ldots,f^{k-1}(C_{j})\}$. In this case, we refer to the change in the portrait that results from applying $\phi$ as “applying Function 11 to $C_{j}$.” The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
Proof.
Notice that $\phi$ exchanges $f^{k-1}(C_{j})$ and $\diamond$, which is a preimage of $f^{2}(C_{j})$ under $f$. Therefore, $f^{k-1}(C_{j})$ now maps to $f^{2}(C_{1})$ (instead of $C_{j}$). Because $\phi$ can be defined to be supported on a neighborhood of an arc between $f^{k-1}(C_{j})$ and $\diamond$, which is disjoint from the other critical and postcritical points, $(f\phi)^{\ell}(C_{j})=f^{\ell}(C_{j})$ for $21\leq\ell\leq k-$ 2. In particular, the pre-period of $f\phi$ is $\{C_{j},f(C_{j})\}$ and the period of $f\phi$ is $\{f^{2}(C_{j}),\ldots,f^{k-1}(C_{j})\}$. Therefore $f\phi$ does not change the number of components of the portrait of $f$. ∎
Function 12 (Create two disjoint pre-periods).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ such that one critical point is in the cycle of the other. Keeping to our conventions, assume that $C_{2}$ is in the cycle of $C_{1}$. Let $r$ be the length of the pre-period of $C_{1}$, let $k$ be the length of the cycle, and suppose that $C_{2}=f^{q}(C_{1})$, where $q>r\geq 3$. Let $\diamond$ be a pre-image of $f^{r-1}(C_{1})$ that is not post-critical. Let $\phi$ be a half twist that swaps $\diamond$ with $f^{q-1}(C_{1})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ is such that $C_{1}$ and $C_{2}$ are both strictly pre-periodic and their pre-periods are disjoint. Moreover, the length of the pre-period of $C_{1}$ equals $r-1$, and the length of the pre-period of $C_{2}$ is $r+k-q+1$, and the length of the cycle is $q-k+1$. The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
[…]
Figure 9. Applying Function 12 changes the pre-period structure of the portrait, but not the number of components.
Proof.
The branched cover $f\phi$ maps $f^{q-1}(C_{1})$ to $f^{r-1}(C_{1})$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc between $\diamond$ and $f^{q-1}(C_{1})$, the maps $f$ and $f\phi$ agree on all critical and post-critical points except for $f^{q-1}(C_{1})$. Then the portrait of $f\phi$ has cycle consisting of $\{f^{r-1}(C_{1}),f^{r}(C_{1}),f^{r+1}(C_{1}),\ldots,f^{q-1}(C_{1})\}$. Moreover, both $C_{1}$ and $C_{2}$ are strictly pre-periodic. The pre-period of $C_{1}$ consists of $\{C_{1},f(C_{1}),\ldots,f^{r-2}(C_{1})\}$, which is of length $r-1$. The pre-period of $C_{2}$ consists of $\{f^{q}(C_{1})=C_{2},f^{q+1}(C_{1}),\ldots,f^{r+k-1}(C_{1})\}$, which is of length $r+k-q$. Notice that the pre-periods of $C_{1}$ and $C_{2}$ are disjoint. ∎
Function 13 (Create distinct and intersecting pre-periods).
Take as input a marked branched cover $f$ with marked set $M$ and associated portrait $\Gamma$ such that one critical point is in the cycle of the other. Keeping to our conventions, assume that $C_{2}$ is in the cycle of $C_{1}$. Let $r$ be the length of the pre-period of $C_{1}$, let $k$ be the length of the cycle, and suppose that $C_{2}=f^{q}(C_{1})$, where $q\geq r+2$. Let $\diamond$ be a pre-image of $f^{r+1}(C_{1})$ that is not post-critical. Let $\phi$ be a half twist that swaps $\diamond$ with $f^{q-1}(C_{1})$. Output $f\phi$, the marked set $M\cup\{\diamond\}$, the portrait of $f\phi$, and the mapping class $\phi$. The portrait of $f\phi$ is such that $C_{1}$ and $C_{2}$ are both strictly pre-periodic and intersect non-trivially and each contains a post-critical point that is unique from the other. The branched cover $f\phi$ does not change the number of portrait components of $\Gamma_{f}$.
[…]
Figure 10. Applying Function 13 changes the pre-period structure of the portrait, but not the number of components.
Proof.
The branched cover $f\phi$ maps $f^{q-1}(C_{1})$ to $f^{r+1}(C_{1})$. Since $\phi$ can be chosen so that it is supported on a neighborhood of an arc between $\diamond$ and $f^{q-1}(C_{1})$, the maps $f$ and $f\phi$ agree on all critical and post-critical points except for $f^{q-1}(C_{1})$. Then the portrait of $f\phi$ has cycle consisting of $\{f^{r+1}(C_{1}),f^{r+2}(C_{1}),\ldots,f^{q-1}(C_{1})\}$. Moreover, both $C_{1}$ and $C_{2}$ are strictly pre-periodic. The pre-period of $C_{1}$ that is disjoint from the pre-period of $C_{2}$ consists of $\{C_{1},f(C_{1}),\ldots,f^{r-1}(C_{1})\}$, which is of length $r$. The pre-period of $C_{2}$ that is disjoint from the pre-period of $C_{1}$ consists of $\{f^{q}(C_{1})=C_{2},f^{q+1}(C_{1}),\ldots,f^{r+k-1}(C_{1})\}$, which is of length $r+k-q$. The intersection of the pre-periods is the point $\{f^{r}(C_{1})\}$. Therefore $f\phi$ does not change the number of components of the portrait $\Gamma$. ∎
Proposition 3.1.
The set of abstract portraits that are compatible with a quadratic branched cover are closed under Functions 1-13.
Proof.
@@
Let $f$ be an abstract portrait compatible with a quadratic branched cover and let $\phi$ be a homeomorphism used in one of Functions 1-13. We claim that $f\phi$ is compatible with a quadratic branched cover. First we note that conditions (1)-(5) in the definition of abstract portrait are used to determine which points in the marked set of $f\phi$ are in the portrait $\Gamma_{f\phi}$ and which points are distinguished as critical points in $\Gamma_{f\phi}$. Because these serve to define the points, they do not need to be verified.
@@
Functions 1, 3, 5, 7, 10, 11, 12, and 13 add a point that is not in $\Gamma$ to the portrait, and therefore do not swap a critical point with the origin of any edge in $\Gamma$. The homeomorphism in Function 2 swaps two points, but at least one is not critical. Conditions (2) and (3) in Function 2 ensure that if the other point is critical, it is not the origin of an edge that has the same terminus as another edge. Functions 4 and 6 swap two non-critical points. In Function 8, the requirement that each component contains at least two points ensures that neither $f(C_{1})$ nor $f(C_{2})$ is critical. Therefore Function 8 swaps two non-critical points. Function 9 has six cases, but all cases either swap two critical points, two non-critical points, or swap a point that is not in $\Gamma$ with a point that is. Therefore no cases of Function 9 swap a critical point with a non-critical origin of an edge in $\Gamma$.
@@ 4. Algorithm
Let $f:S^{2}\rightarrow S^{2}$ be a post-critically finite branched cover with portrait $\Gamma_{f}$. Let $\Gamma$ be a dynamical portrait that is compatible with a quadratic branched cover. Note that $\Gamma$ will contain two points distinguished as critical points $C_{1}$ and $C_{2}$. The algorithm below describes a sequence of portraits $\Gamma^{0},\cdots,\Gamma^{n}$ with $\Gamma_{f}=\Gamma^{0}$ and $\Gamma=\Gamma^{n}$. For each step of the algorithm, we compare the portrait $\Gamma^{i}$ to the portrait $\Gamma^{n}$ and apply one of the functions from Section 3 to obtain the next portrait $\Gamma^{i+1}$. The algorithm also tracks the marked set of the branched cover and the mapping class in each step. Note that the vertices of $\Gamma^{i}$ may be a proper subset of $M_{i}$.
Components of $\Gamma_{f}$ and $\Gamma$. Because a portrait of a quadratic branched cover can have at most two (non-empty) components, we will consider each portrait as a pair of (possibly empty) components $\Gamma_{f}=(\Gamma^{f}_{1},\Gamma^{f}_{2})$ and $\Gamma=(\Gamma_{1},\Gamma_{2})$. If one component is empty, we will always label the empty component as the second entry. Moreover, we will say that a portrait has one component if it has one non-empty component and one empty component. When $\Gamma$ has two (non-empty) components, the choice of labeling $\Gamma_{1}$ and $\Gamma_{2}$ is arbitrary. However, if one component of $\Gamma$ is equal to $\Gamma^{1}_{1}$ (or $\Gamma_{2}^{1}$), then the algorithm may require fewer steps if we choose $\Gamma_{1}$ (respectively $\Gamma_{2}$) to be the component that is equal. It is possible that the algorithm could be improved by choosing $\Gamma_{1}$ efficiently, but we suspect the result will be a minimal improvement.
The algorithm. Steps 4- 4 below give a sequence of pairs of portrait components $\{(\Gamma^{i}_{1},\Gamma^{i}_{2})\}$ so that:
  • $\bullet$
    $\Gamma^{i+1}_{j}$ differs from $\Gamma^{i}_{j}$ by a function from Section 3,
  • $\bullet$
    there exists $n\geq 1$, so that $\Gamma^{n}_{j}=\Gamma_{j}$, and
  • $\bullet$
    the critical points of $f_{n}$ are the points distinguished as critical points for $\Gamma$.
We initialize the algorithm with the pair $(\Gamma^{0},f_{0})=(\Gamma_{f},f)$, the target portrait $\Gamma$, the set $M_{0}$ consisting of $P_{f}$ and the critical points of $\Gamma_{f}$, the identity mapping class $\phi$, and the sequence (at this stage consisting of the single entry) $\{(\Gamma^{0},f_{0})\}$.
@@
Take as input the pair $(\Gamma^{i},f_{i})$, the portrait $\Gamma$, the marked set $M_{i}$, a mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$. As above, each $\Gamma^{i}$ is a pair of (possibly empty) components.
@@
Compare the number of components of $\Gamma^{i}$ and $\Gamma$. If both $\Gamma^{i}$ and $\Gamma$ have two non-empty components, return $(\Gamma^{i},f_{i})$, the portrait $\Gamma$, the marked set $M_{i}$, the mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$, and continue to Step 4. If $\Gamma$ and $\Gamma^{i}$ both have one component, return $(\Gamma^{i},f_{i})$, the portrait $\Gamma$, the marked set $M_{i}$, the mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$ and continue to Step 4.
@@
If $\Gamma^{i}$ has one component and $\Gamma$ has two components, then we apply Function 9. In particular, if the component of $\Gamma^{i}$ is a cycle, apply Function 9 (1). If the component of $\Gamma^{i}$ has one pre-period, apply Function 9 (2). If $\Gamma^{i}$ has two pre-periods where one is contained in the other, apply Function 9 (3). If $\Gamma^{i}$ is a portrait which has two disjoint pre-periods, apply Function 9 (4). If $\Gamma^{i}$ is a portrait which has two pre-periods which intersect non-trivially and each contains a point that is distinct from the other, apply Function 9 (5). Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 9. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If $\Gamma^{i}$ has two components, $\Gamma$ has one component, and one of the components of $\Gamma^{i}$ has exactly one vertex, apply Function 1. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If $\Gamma^{i}$ has two components, $\Gamma$ has one component, and both components have at least two vertices, apply Function 8. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 8. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
Step 2: Pre-period structure starting with two components.
@@
Take as input the pair $(\Gamma^{i},f_{i})$, the portrait $\Gamma$, the marked set $M_{i}$, a mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$. Let $\Gamma^{i}=(\Gamma_{1}^{i},\Gamma_{2}^{i})$ and $\Gamma=(\Gamma_{1},\Gamma_{2})$. We may assume that $\Gamma^{i}$ (and hence $\Gamma$) has two non-empty components because Step 4 is only called in this case. Let $C_{1}^{i}$ and $C_{2}^{i}$ be the critical points of $f_{i}$ and, as above, let $C_{1}$ and $C_{2}$ be the points distinguished as critical points in $\Gamma$.
If both components of $\Gamma$ are cycles and both components of $\Gamma^{i}$ are cycles, continue to Step 4. If both components of $\Gamma$ are strictly pre-periodic and both components of $\Gamma^{i}$ are strictly pre-periodic, continue to Step 4. If one component of $\Gamma$ is a cycle and the other has a pre-period and one component of $\Gamma^{i}$ is a cycle and the other has a pre-period, label the component of $\Gamma$ that is a cycle by $\Gamma_{1}$, and label the component of $\Gamma^{i}$ that is a cycle by $\Gamma^{i}_{1}$. Label the component of $\Gamma$ that is strictly pre-periodic by $\Gamma_{2}$, and label the component of $\Gamma^{i}$ that is strictly pre-periodic by $\Gamma^{i}_{2}$. Continue to Step 4.
Otherwise, either $\Gamma_{1}^{i}$ and $\Gamma_{1}$ have a different pre-period structure or $\Gamma_{2}^{i}$ and $\Gamma_{2}$ have a different pre-period structure (or both).
If $C_{1}^{i}$ is strictly pre-periodic under $f_{i}$ and $C_{1}$ is in a cycle of $\Gamma_{1}$, then apply Function 7 to $C_{1}^{i}$. Notice that we can immediately call Function 7 since $f(C_{1})\neq C_{2}$ by default since this step intakes a portrait with two non-empty components. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 7. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If $C_{1}^{i}$ is in the cycle of $\Gamma^{i}_{1}$ and $C_{1}$ is not in the cycle of $\Gamma_{1}$:
{adjustwidth}
1cm       If the cycle of $\Gamma^{1}_{j}$ has length less than 3, apply Function 1 to $\Gamma^{i}_{1}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If the cycle of $\Gamma^{i}$ has length at least 3, apply Function 11 to $C_{1}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 11. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
@@
If $C_{2}^{i}$ is strictly pre-periodic under $f_{i}$ and $C_{2}$ is in a cycle of $\Gamma_{1}$, then apply Function 7 to $C_{2}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 7. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If $C_{2}^{i}$ is in the cycle of $\Gamma^{i}_{2}$ and $C_{2}$ is not in the cycle of $\Gamma_{2}$:
{adjustwidth}
1cm       If the cycle of $\Gamma^{i}_{2}$ has length less than 3, apply Function 1 to $\Gamma^{i}_{2}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
If the cycle of $\Gamma^{i}$ has length at least 3, apply Function 11 to $C_{2}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 11. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
Step 3: Pre-period structure starting with one component.
Take as input the pair $(\Gamma^{i},f_{i})$, the portrait $\Gamma$, the marked set $M_{i}$, a mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$. Let $\Gamma^{i}=(\Gamma_{1}^{i},\Gamma_{2}^{i})$ and $\Gamma=(\Gamma_{1},\Gamma_{2})$. We may assume that $\Gamma^{i}_{1},\Gamma_{1}\neq\emptyset$ and $\Gamma^{i}_{2},\Gamma_{2}=\emptyset$.
If $\Gamma_{1}^{i}$ and $\Gamma_{1}$ both have no pre-periods (both critical points are periodic), then continue to Step 4. If $\Gamma_{1}^{i}$ and $\Gamma_{1}$ both have one pre-period (one critical point is strictly pre-periodic and the other is periodic), then continue to Step 4.
If $\Gamma_{1}^{i}$ and $\Gamma_{1}$ both have two pre-periods, we compare the pre-period structure of $\Gamma^{i}_{1}$ to the pre-period structure of $\Gamma_{1}$. If $\Gamma^{i}_{1}$ and $\Gamma_{1}$ both have the same pre-period structure, that is both have:
  • $\bullet$
    two disjoint pre-periods,
  • $\bullet$
    two pre-periods that are distinct and intersecting, or
  • $\bullet$
    two pre-periods in which one is contained in the other,
then continue to Step 4.
In the remaining cases, $\Gamma_{1}^{i}$ and $\Gamma_{1}$ either have a different number of pre-periods or a different pre-period structure, which we reconcile in two steps. First, if $\Gamma_{1}^{i}$ has zero or two pre-periods, apply a function to obtain a portrait with one pre-period. Second, once $\Gamma_{1}^{i}$ has one pre-period, follow the appropriate case to obtain a portrait with the same pre-period structure as $\Gamma$.
If $\Gamma^{i}_{1}$ is a cycle (that is, both $C_{1}^{i}$ and $C_{2}^{i}$ are periodic) and $\Gamma_{1}$ is not:
{adjustwidth}
1cm
If $f(C_{1}^{i})=C_{2}$, apply Function 1 to $C_{2}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Repeat Step 4.
If $f(C_{2}^{i})=C_{1}$, apply Function 1 to $C_{1}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Repeat Step 4.

### v2

@@ 3. Portrait functions
Let $f$ be an abstract portrait compatible with a quadratic branched cover and let $\phi$ be a homeomorphism used in one of Functions 1-3. We claim that $f\phi$ is compatible with a quadratic branched cover. First we note that conditions (1)-(5) in the definition of abstract portrait are used to determine which points in the marked set of $f\phi$ are in the portrait $\Gamma_{f\phi}$ and which points are distinguished as critical points in $\Gamma_{f\phi}$. Because these serve to define the points, they do not need to be verified.
@@
Function 1 has six cases, but all cases either swap two critical points, two non-critical points, or swap a point that is not a vertex of $\Gamma$ with a point that is a vertex of $\Gamma$. Therefore no cases of Function 1 swap a critical point with a non-critical origin of an edge in $\Gamma$. Function 2 adds a point that is not in $\Gamma$ to the portrait, and therefore do not swap a critical point with the origin of any edge in $\Gamma$. The homeomorphism in Function 3 swaps two post-critical points, but is only applied under the assumption that the components are both cycles and therefore has no points that are the terminus of two edges. ∎
@@ 4. Algorithm
Let $g,h:S^{2}\rightarrow S^{2}$ be a post-critically finite branched covers with portrait $\Gamma_{g}$ and $\Gamma_{h}$, respectively. Let $\Gamma$ denote the portrait of $p(z)=z^{2}$, which consists of the two fixed critical points 0 and $\infty$. The algorithm below describes a sequence of portraits $\Gamma^{0},\cdots,\Gamma^{n}$ with $\Gamma_{g}=\Gamma^{0}$ and $\Gamma=\Gamma^{n}$. For each step of the algorithm, we compare the portrait $\Gamma^{i}$ to the portrait $\Gamma$ and apply one of the functions from Section 3 to obtain the next portrait $\Gamma^{i+1}$. Then we repeat the algorithm for $h$. The algorithm tracks the marked set in each step and the mapping class that is composed in each step. Note that in each step, the domain is a sphere, but the marked set may be a proper superset of the marked set in the previous step. We are able to compose mapping classes from each step, possibly by applying capping homomorphisms to mapping classes from some steps [4]. The final marked set $M$ is the union of the marked set in each step of the algorithm for both $g$ and $h$. The final mapping class $\phi$ is the composition of the mapping classes from each step of the algorithm for $g$ and the inverse of the mapping class from each step of the algorithm for $h$.
Components of $\Gamma_{g}$, $\Gamma_{h}$, and $\Gamma$. Because a portrait of a quadratic branched cover can have at most two (non-empty) components, we will consider each portrait as a pair of (possibly empty) components $\Gamma_{g}=(\Gamma^{g}_{1},\Gamma^{g}_{2})$, $\Gamma_{h}=(\Gamma^{h}_{1},\Gamma^{h}_{2})$, and $\Gamma=(\Gamma_{1},\Gamma_{2})$. If one component is empty, we will always label the empty component as the second entry. Moreover, we will say that a portrait has one component if it has one non-empty component and one empty component.
The algorithm. Steps 4- 4 below give a sequence of portraits $\Gamma^{i}=\{(\Gamma^{i}_{1},\Gamma^{i}_{2})\}$ so that $\Gamma^{i+1}_{j}$ differs from $\Gamma^{i}_{j}$ by a function from Section 3 and there exists $n\geq 1$, so that $\Gamma^{n}=\Gamma$.
We first initialize the algorithm with the pair $(\Gamma^{0},f_{0})=(\Gamma_{g},g)$, the set $M_{0}$ consisting of $P_{g}$ and the critical points of $g$, the identity mapping class $\phi$, and the sequence (at this stage consisting of the single entry) $\{(\Gamma^{0},f_{0})\}$. After completing the algorithm for $g$, we initialize the algorithm with the pair $(\Gamma^{0},f_{0})=(\Gamma_{h},h)$, the set $M_{0}$ consisting of $P_{h}$ and the critical points of $h$, the identity mapping class $\phi$, and the (single entry) sequence $\{(\Gamma_{0},f_{0})\}$.
@@
Take as input the pair $(\Gamma^{i},f_{i})$, the marked set $M_{i}$, a mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$. As above, each $\Gamma^{i}$ is a pair of (possibly empty) components.
@@
If $\Gamma^{i}$ has two non-empty components, return $(\Gamma^{i},f_{i})$, the marked set $M_{i}$, the mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$, and continue to Step 4.
@@
If $\Gamma^{i}$ has one component, then we apply Function 1. In particular, if the component of $\Gamma^{i}$ is a cycle, apply Function 1 (1). If the component of $\Gamma^{i}$ has one pre-period, apply Function 1 (2). If $\Gamma^{i}$ has two pre-periods where one is contained in the other, apply Function 1 (3). If $\Gamma^{i}$ is a portrait which has two disjoint pre-periods, apply Function 1 (4). If $\Gamma^{i}$ is a portrait which has two pre-periods which are distincting and intersecting, apply Function 1 (5). Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 1. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.
Step 2: Pre-period structure.
@@
Take as input the pair $(\Gamma^{i},f_{i})$, the marked set $M_{i}$, a mapping class $\phi_{i}$, and the sequence $\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\}$. Let $\Gamma^{i}=(\Gamma_{1}^{i},\Gamma_{2}^{i})$ and $\Gamma=(\Gamma_{1},\Gamma_{2})$. We may assume that $\Gamma^{i}$ has two non-empty components because Step 4 is only called in this case. Let $C_{1}^{i}$ and $C_{2}^{i}$ be the critical points of $f_{i}$ and, as above.
If both components of $\Gamma^{i}$ are cycles, continue to Step 4.
@@
If $C_{1}^{i}$ is strictly pre-periodic under $f_{i}$, then apply Function 2 to $C_{1}^{i}$. Let $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i+1})$ be the output of Function 2. Return $(f_{i+1},M_{i+1},\Gamma_{i+1},\phi_{i}\circ\phi_{i+1},\{(\Gamma^{0},f_{0}),\cdots,(\Gamma^{i+1},f_{i+1})\})$. Continue to Step 4.